# On a generalisation of Cameron’s base size conjecture

### v2

@@ Abstract.
In this paper, we complete the proof of a natural generalisation of Cameron’s conjecture. Our main result states that if $G$ is an almost simple group and $H_{1},\ldots,H_{k}$ are any core-free non-standard maximal subgroups of $G$ with $k\geqslant 7$, then $G$ has a regular orbit on $G/H_{1}\times\cdots\times G/H_{k}$, noting that Cameron’s original conjecture corresponds to the special case where the $H_{i}$ are pairwise conjugate subgroups. In addition, we show that the same conclusion holds with $k\geqslant 6$, unless $G={\rm M}_{24}$ and each $H_{i}$ is isomorphic to ${\rm M}_{23}$. For example, this means that if $G$ is a simple exceptional group of Lie type and $H_{1},\ldots,H_{6}$ are proper subgroups of $G$, then there exist elements $g_{i}\in G$ such that $\bigcap_{i}H_{i}^{g_{i}}=1$. By applying recent work in a joint paper with Burness, we may assume $G$ is a group of Lie type and our proof uses probabilistic methods based on fixed point ratio estimates.
@@ 1. Introduction
Let $G\leqslant\operatorname{Sym}(\Omega)$ be a transitive permutation group on a finite set $\Omega$. Recall that a base for $G$ is a subset of $\Omega$ whose pointwise stabiliser is trivial. We write $b(G,\Omega)$ for the base size of $G$, which is the minimal cardinality of a base for $G$. This classical invariant has been extensively studied since the early days of group theory in the nineteenth century, finding a wide range of applications to other areas of mathematics such as graph theory and model theory, for example. Bases are also fundamental in the computational study of finite groups and the implementation of efficient algorithms. There is an extensive literature on bases for permutation groups and we refer the reader to the survey articles [7, 62, 67] and [20, Section 5] for further details.
Determining the base size of a permutation group is a fundamental problem in group theory and it has attracted significant interest in recent years. In particular, there has been particular focus in the study of base sizes of primitive groups, motivated by some highly influential conjectures by Babai, Cameron, Kantor, and Pyber stated in the 1990s (Recall that a transitive permutation group $G\leqslant{\rm Sym}(\Omega)$ is primitive if a point stabiliser is a maximal subgroup of $G$). All those conjectures have now been resolved.
Recall that a finite group $G$ is almost simple if there exists a nonabelian simple group $G_{0}$ such that $G_{0}\trianglelefteqslant G\leqslant{\rm Aut}(G_{0})$. We call $G_{0}$ the socle of $G$. We say that a subgroup $H$ of a finite group $G$ is core-free if $\bigcap_{g\in G}H^{g}=1$. Note that if $G$ is almost simple, then $H\leqslant G$ is core-free if it does not contain $G_{0}$. Recall the following definition:
@@
We will define subspace subgroups of classical groups precisely in Section 3 (see Definition 3.4), but roughly speaking, if $G$ is a classical group, then $H$ is standard if it is reducible. That is, $H\cap G_{0}$ acts reducibly on the natural module. Note that if there exists an isomorphism of $G$ to some other group that maps $H$ to a standard subgroup, then we will also say that $H$ is standard in this case. For instance, we note that ${\rm PSL}_{4}(2)\cong A_{8}$ and the subgroup of type ${\rm GL}_{2}(4)$ gets mapped to the intransitive subgroup $(S_{3}\times S_{5})\cap A_{8}$ by an isomorphism. We will therefore say that the subgroup of type ${\rm GL}_{2}(4)$ is standard in ${\rm PSL}_{4}(2)$. Those cases are handled in Table 3.
If $G\leqslant{\rm Sym}(\Omega)$ is a primitive group with point stabiliser $H$, then we will say that the action of $G$ on $G/H$ is non-standard if $H$ is non-standard. It is natural to distinguish between standard and non-standard actions when studying bases. More specifically, we observe that there is a dichotomy between standard and non-standard actions in terms of base sizes. In particular, standard actions tend to admit ‘large bases’. For example, the base sizes of alternating and symmetric groups in standard actions have been determined due to work by several authors (see [21, 35, 45, 46, 68, 69]) and we observe that in the majority of cases, $b(G,\Omega)$ grows with $n$, and consequently $b(G,\Omega)$ can get arbitrarily large as $n\rightarrow\infty$. As we now explain, almost simple groups in non-standard actions exhibit a completely different behaviour with respect to base sizes.
It was conjectured by Cameron and Kantor in [29] that if $G$ is a primitive almost simple permutation group in a non-standard action, then $b(G,\Omega)$ is bounded above by an absolute constant independent of $G$. This was proved by Liebeck and Shalev in [55]. In response, Cameron conjectured in [28, p.122] that the best possible value of this absolute constant is $7$. This conjecture was proved in a series of papers by Burness et al. [18, 23, 25, 26]. In particular, we have the following result:
Theorem.
Let $G\leqslant{\rm Sym}(\Omega)$ be an almost simple primitive permutation group with non-standard point stabiliser $H$. Then $b(G,\Omega)\leqslant 7$ with equality if and only if $(G,H)=({\rm M}_{24},{\rm M}_{23})$.
The aim of this paper is to generalise Theorem Theorem in the setting we describe below:
Let $G$ be a finite group and let $H_{1},\ldots,H_{k}$ be a collection of core-free subgroups of $G$, allowing repetitions. Consider the natural action of $G$ on the Cartesian product
@@
for some elements $g_{i}\in G$. Following [2], we say that a $k$-tuple $(H_{1},\ldots,H_{k})$ of core-free subgroups of $G$ is regular if $G$ has a regular orbit on $X$, and non-regular otherwise. We will always be interested in the regularity of subgroup tuples all of whose components are core-free, and so, given a subgroup tuple, we will always assume that its components are core-free. We will also say that $\tau$ is conjugate if the $H_{i}$ are pairwise conjugate. Finally, we call the minimal $r$ such that all $r$-tuples of core-free subgroups of $G$ are regular the regularity number of $G$, which we denote by $R(G)$.
Recall that a transitive group $G\leqslant{\rm Sym}(\Omega)$ has a base of size $k$ if and only if $G$ has a regular orbit on $\Omega^{k}$. Therefore, a conjugate tuple $(H_{1},\ldots,H_{k})$ is regular if and only if $G$ admits a base of size $k$ in its natural action on $G/H_{1}$. This implies that $k\geqslant b(G,H_{1})$.
If $G$ is almost simple and all the components of a tuple $\tau=(H_{1},\ldots,H_{k})$ are non-standard, we will say that $\tau$ is non-standard. Hence, in this language, Theorem Theorem asserts that every conjugate non-standard $7$-tuple of $G$ is regular.
The regularity of subgroup tuples was introduced and studied in [2] for almost simple groups with alternating and sporadic socle, where Burness and the author stated generalisations of some of the most influential base size conjectures in the regularity setting, including Cameron’s conjecture. In particular, if $G$ is a finite almost simple group and $R_{\rm ns}(G)$ denotes the minimal integer $r$ such that all non-standard tuples of $G$ are regular, then they conjecture that $R_{\rm ns}(G)\leqslant 7$, with equality if and only if $G={\rm M}_{24}$. Note that this conjecture is a vast generalisation of Theorem Theorem. As noted above, Theorem Theorem implies that all conjugate non-standard $7$-tuples of a finite almost simple group $G$ are regular and it is claimed that the conjugacy assumption can be removed.
The above conjecture was proved in [2] for almost simple groups with alternating and sporadic socle, which reduces the problem to the almost simple groups of Lie type. The main aim of this paper is to complete the proof of this conjecture. We now state our main theorem:
@@
Let $G$ be a finite almost simple group. Then $R_{\rm ns}(G)\leqslant 7$ with equality if and only if $G={\rm M}_{24}$.
It is shown in [2], that up to conjugacy, the only non-regular non-standard $6$-tuple when $G={\rm M}_{24}$ is $(H,H,H,H,H,H)$, where $H={\rm M}_{23}$. Moreover, note that if $G$ is sporadic or exceptional, then all of its subgroups are non-standard, so Theorem 1 asserts that $R(G)\leqslant 7$ in this case.
As we will see in Section 3, we can prove a stronger statement than Theorem 1 for classical groups. In particular, if $G$ is an almost simple classical group, then we will show that every non-standard $4$-tuple is regular, unless $G={\rm U}_{6}(2).2$, in which case $R_{\rm ns}(G)=5$ and the only non-regular non-standard $4$-tuple up to conjugacy is $(H,H,H,H)$ with $H={\rm U}_{4}(3).2^{2}$. Combining this with the results in [2], we obtain a more refined version of Theorem 1, which we state below:
@@
    If $G_{0}$ is an exceptional group, then $R_{\rm ns}(G)=R(G)\leqslant 6$.
We now briefly discuss the methods we use to prove Theorem 1 for groups of Lie type. These are similar to those used in the proof of Cameron’s conjecture. In particular, our main tool will be a generalised version of the powerful probabilistic method based on fixed point ratio estimates that was first introduced in [63] by Liebeck and Shalev for the study of bases.
Let $G$ be a finite group and let $\tau=(H_{1},\ldots,H_{k})$ be a tuple of core-free subgroups of $G$. We will write $X_{\tau}=G/H_{1}\times\cdots\times G/H_{k}$ and $\mathbb{P}(G,\tau)$ for the probability that a randomly chosen point in $X_{\tau}$ is in a regular orbit. Our general approach will be to show that
@@
Note that generally, it is impossible to compute $Q(G,\tau)$ for a finite group $G$. However, in most cases it is sufficient to work with the upper bound below:
@@
where $x_{1},\ldots,x_{t}$ is a set of representatives of the prime order conjugacy classes in $G$ and ${\rm fpr}(x,G/H)$ denotes the fixed point ratio of $x$ in the natural action of $G$ on the cosets of $H$. That is, the proportion of fixed points of $x$ in this action. We refer the reader to Section 2.1 for more information.
If $G$ is an almost simple classical group with socle $G_{0}$ with natural module $V$ and $\tau$ is a non-standard tuple of $G$, then our main tool when estimating $\widehat{Q}(G,\tau)$ the following ‘zeta type function’
@@
where $\mathscr{C}$ denotes the set of prime order $G_{0}$-conjugacy classes of $G$. More specifically, Burness obtains upper bounds on ${\rm fpr}(x,G/H)$ for a prime order element $x$ and a non-standard subgroup $H$ in terms of the size of the $G$-class of $x$ in [14, 15, 16, 17]. Using these allows us to bound $\widehat{Q}(G,\tau)$ by $\eta_{G}(t)$ for some $t\in(0,1)$, and our problem reduces to showing that $\eta_{G}(t)<1$. In some cases, we will need stronger estimates on ${\rm fpr}(x,G/H)$ than the ones obtained [14, 15, 16, 17], and we obtain those in Section 4. Finally, when ${\rm dim}V\leqslant 5$, using $\eta_{G}(t)$ will no longer be possible, so we take a more direct approach to estimate $\widehat{Q}(G,\tau)$.
If $G$ is exceptional, then we obtain a character theoretic approach instead. In particular, if $H$ is a maximal subgroup of $G$ and $\Omega=G/H$, then we have
${\rm fpr}(x,G/H)=\frac{\chi(x)}{|\Omega|},$
where $\chi(x)$ is the corresponding permutation character. Noting this, we can employ powerful tools developed by Lawther, Liebeck and Seitz in [51] to estimate ${\rm fpr}(x,G/H)$. In this way, for every prime order element $x\in G$, we are able to compute
@@
where $\mathcal{M}_{G}$ denotes the set of maximal subgroups of $G$. This allows us to obtain a uniform upper bound on $\widehat{Q}(G,\tau)$ that applies to any $6$-tuple $\tau$ of maximal subgroups of $G$.
The main challenge when the conjugacy assumption on our subgroup tuples is removed is that we need to adopt a uniform approach that is applicable to any non-standard tuple. In particular, our arguments no longer depend on a particular action of our group $G$, which means that we are always bound to worst case fixed point ratio estimates over all maximal subgroups of $G$ for all prime order elements of $G$.
Existing worst case bounds in the literature for a given element $x\in G$ might be good enough for a particular action of $G$, but might not be sufficient in this general setting. This means that we will often need to obtain stronger fixed point ratio estimates for our purposes. We state those in Theorem 4.1 for the classical groups. For exceptional groups, we record worst case fixed point ratio estimates over non-parabolic actions in Propositions 8.1, 8.2, 8.3, and 8.4. Our main source for obtaining those is inspection of [25] and [19], and in some instances we need to improve what currently is available. We give detailed information in Sections 2, 4 and 8.
@@
The heavily probabilistic approach we adopt to prove Theorem 1 allows us to also obtain an asymptotic result. In particular, if $G$ is an almost simple group of Lie type, $\mathcal{P}$ denotes the set of non-standard subgroups of $G$ and
$\mathbb{P}(G,c)=\min\{\mathbb{P}(G,\tau)\,:\,\tau\in\mathcal{P}^{c}\},$
then we have the following theorem:
@@
This paper is organised as follows: In Section 2 we record some preliminary results. In Section 3, we proceed by recording some preliminaries that only apply to classical groups. We first discuss subgroup structure and conjugacy classes of classical groups, and then proceed to give a precise definition of a subspace subgroup of a classical group. The rest of the section is devoted to probabilistic and computational methods, which are the key tools that we will be using.
@@
Section 4 is devoted to improving existing fixed point ratio estimates for low-dimensional classical groups, in Section 5 we prove some results relating to the ‘zeta-type’ function introduced in Definition 3.10, and Section 6, the final section on classical groups is devoted to the proof of Theorem 3.1, a refinement of Theorem 1 for classical groups.
@@
In Section 7 we record some preliminaries that only apply to exceptional groups. In particular, the section is divided into two subsections, where we briefly discuss subgroup structure and conjugacy classes respectively. In Section 8 we record fixed point ratio estimates for non-parabolic actions of exceptional groups, and then in Section 9 we proceed to discuss fixed point ratios for parabolic actions. Finally, Sections 10.1, 10.2, and 10.3 are devoted to the proof of Theorem 7.1, a refinement of Theorem 1 for exceptional groups.
@@
Notation. Our notation is fairly standard. We write ${\rm Inndiag}(G_{0})$ for the group of inner-diagonal automorphisms of a simple group of Lie type $G_{0}$. Moreover, we will also often write $\pi(n)$ to denote the set of prime divisors of a positive integer $n$ and $\pi_{a}(n)$ to denote the set of prime divisors of $n$ that are no less than $a$. In a similar fashion, we will denote the set of prime numbers at most $s$ by $\pi_{s}$ and the set of odd prime numbers at most $s$ by $\pi^{\prime}_{s}$. For a group $G$, we will denote the number of elements of order $s$ in $G$ by $i_{s}(G)$. For classical groups, we write ${\rm PSL}^{+}_{n}(q)={\rm PSL}_{n}(q)={\rm L}_{n}(q)$ and ${\rm PSL}^{-}_{n}(q)={\rm PSU}_{n}(q)={\rm U}_{n}(q)$, We use the same notation and terminology as [22] for field, graph, and graph-field automorphisms. Finally, we follow the definition of [47] for the $\mathscr{C}_{i}$ and $\mathscr{S}$ classes of maximal subgroups of classical groups. We will also be using the standard notation $E_{6}^{\epsilon}(q)$ for the finite exceptional groups of type $E_{6}$, where $E_{6}^{+}(q)=E_{6}(q)$ and $E_{6}^{-}(q)={}^{2}E_{6}(q)$. We will introduce more technical terminology, when and where appropriate in the paper.

### v3

# A generalisation of Cameron’s base size conjecture

@@ Abstract.
In this paper, we complete the proof of a natural generalisation of Cameron’s conjecture. Our main result states that if $G$ is an almost simple group and $H_{1},\ldots,H_{k}$ are any non-standard maximal subgroups of $G$ with $k\geqslant 7$, then $G$ has a regular orbit on $G/H_{1}\times\cdots\times G/H_{k}$, noting that Cameron’s original conjecture corresponds to the special case where the $H_{i}$ are pairwise conjugate subgroups. In addition, we show that the same conclusion holds with $k=6$, unless $G={\rm M}_{24}$ and each $H_{i}$ is isomorphic to ${\rm M}_{23}$. For example, this means that if $G$ is a simple exceptional group of Lie type and $H_{1},\ldots,H_{6}$ are proper subgroups of $G$, then there exist elements $g_{i}\in G$ such that $\bigcap_{i}H_{i}^{g_{i}}=1$. By applying recent work in a joint paper with Burness, we may assume $G$ is a group of Lie type and our proof uses probabilistic methods based on fixed point ratio estimates.
@@ 1. Introduction
Let $G\leqslant\operatorname{Sym}(\Omega)$ be a transitive permutation group on a finite set $\Omega$. Recall that a base for $G$ is a subset of $\Omega$ whose pointwise stabiliser is trivial. We write $b(G,\Omega)$ for the base size of $G$, which is the minimal cardinality of a base for $G$. This classical invariant has been extensively studied since the early days of group theory in the nineteenth century, finding a wide range of applications to other areas of mathematics such as graph theory and model theory, for example. Bases are also fundamental in the computational study of finite groups and the implementation of efficient algorithms (see [72]). There is an extensive literature on bases for permutation groups and we refer the reader to the survey articles [7, 63, 68] and [20, Section 5] for further details.
Determining the base size of a permutation group is a fundamental problem in group theory and it has attracted significant interest in recent years. In particular, there has been a strong focus on studying base sizes of primitive groups, motivated by some highly influential conjectures due to Babai, Cameron, Kantor, and Pyber from the 1990s. (Recall that a transitive permutation group $G\leqslant{\rm Sym}(\Omega)$ is primitive if a point stabiliser is a maximal subgroup of $G$). Although these conjectures have now been resolved, the extensive work and newly introduced methods required to solve those problems has created lasting momentum in the area and several new directions of research have emerged.
A finite group $G$ is almost simple if there exists a nonabelian simple group $G_{0}$ such that $G_{0}\trianglelefteqslant G\leqslant{\rm Aut}(G_{0})$, and we call $G_{0}$ the socle of $G$ (this is the unique minimal normal subgroup of $G$). The possibilities for $G_{0}$ are determined by the Classification of Finite Simple Groups. In particular, $G_{0}$ is either alternating, or a group of Lie type, or one of the so-called sporadic simple groups. These include the family of Mathieu groups ${\rm M}_{11},{\rm M}_{12},{\rm M}_{22},{\rm M}_{23},{\rm M}_{24}$.
We say that a subgroup $H$ of a finite group $G$ is core-free if $\bigcap_{g\in G}H^{g}=1$. Note that if $G$ is almost simple, then $H$ is core-free if and only if it does not contain $G_{0}$. Let us now recall the following key definition:
@@
We will formally define a subspace subgroup of a classical group in Section 3 (see Definition 3.4). But roughly speaking, if $G$ is a classical group, then $H$ is standard if it is reducible, in the sense that $H\cap G_{0}$ acts reducibly on the natural module for $G_{0}$. Note that if there exists an isomorphism of $G$ to some other group $K$ that maps $H$ to a standard subgroup of $K$, then we will also say that $H$ is standard in this case. For instance, there exists an isomorphism $\phi:G\rightarrow K$ with $G={\rm PSL}_{4}(2)$ and $K=A_{8}$, which maps the maximal subgroup $G={\rm GL}_{2}(4).2$ of $G$ to the stabiliser in $K$ of some $3$-element subset of $\{1,\ldots,8\}$, so we will view $H$ as a standard subgroup of $G$. Those cases are handled in Table 3.
If $G$ is an almost simple group in a transitive action with point stabiliser $H$, then we will say that the action of $G$ is non-standard if $H$ is non-standard. It turns out that groups in standard actions behave vastly differently to those in non-standard actions in terms of bases, so it is natural to distinguish between standard and non-standard actions in this setting. In particular, standard actions tend to admit much larger bases. For example, the base sizes of $S_{n}$ and $A_{n}$ in standard actions have been determined due to work by several authors (see [21, 36, 46, 47, 70, 71]) and we observe that in the majority of cases, $b(G,\Omega)$ can get arbitrarily large as $n\rightarrow\infty$. For example, $b(S_{n},\{1,\ldots,n\})=n-1$. On the contrary, as we will now explain, bases of almost simple groups in non-standard actions exhibit a very different behaviour.
It was conjectured by Cameron and Kantor in [30] that if $G$ is a primitive almost simple permutation group in a non-standard action, then $b(G,\Omega)$ is bounded above by an absolute constant independent of $G$. This was proved by Liebeck and Shalev in [56] with an undetermined constant. In response, Cameron conjectured in [29, p.122] that the best possible value of this absolute constant is $7$. This conjecture was proved in a series of papers by Burness et al. [18, 24, 26, 27]. In particular, we have the following result:
Theorem 1.
Let $G\leqslant{\rm Sym}(\Omega)$ be an almost simple primitive permutation group with non-standard point stabiliser $H$. Then $b(G,\Omega)\leqslant 7$, with equality if and only if $(G,H)=({\rm M}_{24},{\rm M}_{23})$.
The aim of this paper is to generalise Theorem 1 in the setting we describe below.
Let $G$ be a finite group and let $H_{1},\ldots,H_{k}$ be a collection of core-free subgroups of $G$, allowing repetitions. Consider the natural componentwise action of $G$ on the Cartesian product
@@
for some elements $g_{i}\in G$. Following [2], we say that a $k$-tuple $\tau=(H_{1},\ldots,H_{k})$ of core-free subgroups of $G$ is regular if $G$ has a regular orbit on $X$, and non-regular otherwise. Here, we are only interested in the regularity of subgroup tuples all of whose components are core-free, and so, given a subgroup tuple, we will always assume that its components are core-free. We will also say that $\tau$ is conjugate if the $H_{i}$ are pairwise conjugate. Finally, we call the minimal $r$ such that every $r$-tuple of core-free subgroups of $G$ is regular the regularity number of $G$, which we denote by $R(G)$.
Observe that a transitive group $G\leqslant{\rm Sym}(\Omega)$ has a base of size $k$ if and only if $G$ has a regular orbit on $\Omega^{k}$. Therefore, a conjugate tuple $(H_{1},\ldots,H_{k})$ is regular if and only if $G$ admits a base of size $k$ in its natural action on $G/H_{1}$, which means that $k\geqslant b(G,H_{1})$.
If $G$ is almost simple and all the components of a tuple $\tau=(H_{1},\ldots,H_{k})$ are non-standard, then we will say that $\tau$ is non-standard. Hence, in this language, Theorem 1 asserts that every conjugate non-standard $7$-tuple of $G$ is regular.
The regularity of subgroup tuples was introduced and studied in [2] for almost simple groups with alternating and sporadic socle, where Burness and the author proposed generalisations of some of the most influential base size conjectures in the regularity setting, including Cameron’s conjecture. In particular, if $G$ is a finite almost simple group and $R_{\rm ns}(G)$ denotes the minimal integer $r$ such that all non-standard tuples of $G$ are regular, then they conjecture that $R_{\rm ns}(G)\leqslant 7$, with equality if and only if $G={\rm M}_{24}$. As noted above, Theorem 1 implies that all conjugate non-standard $7$-tuples of a finite almost simple group $G$ are regular and we are claiming that the same conclusion holds without the restrictive conjugacy condition.
This conjecture was proved in [2] for almost simple groups with alternating and sporadic socle, which reduces the problem to the almost simple groups of Lie type. The main aim of this paper is to complete the proof of this conjecture. We now state our main theorem:
@@
Let $G$ be a finite almost simple group. Then $R_{\rm ns}(G)\leqslant 7$, with equality if and only if $G={\rm M}_{24}$.
It is shown in [2], that up to conjugacy, the only non-regular non-standard $6$-tuple when $G={\rm M}_{24}$ is $(H,H,H,H,H,H)$, where $H={\rm M}_{23}$. Moreover, note that if $G$ is sporadic or exceptional, then every core-free maximal subgroup is non-standard, so Theorem 2 asserts that $R(G)\leqslant 7$ (with equality if and only if $G={\rm M}_{24}$) in these cases.
As we will see in Section 3, we will prove a stronger version of Theorem 2 for classical groups. In particular, if $G$ is an almost simple classical group, then we show that $R_{\rm ns}(G)\leqslant 5$ with equality if and only if $G={\rm U}_{6}(2).2$, in which case the only non-regular $4$-tuple is $(H,H,H,H)$ with $H={\rm U}_{4}(3).2^{2}$. Combining this with [2, Theorem 1(ii) and Theorem 2(i)], we obtain the following more refined version of Theorem 2.
@@
    If $G_{0}$ is an exceptional group, then $R_{\rm ns}(G)=R(G)\leqslant 6$, with equality if $G_{0}=E_{6}(q)$ or $E_{7}(q)$.
We note that it is still an open problem to determine all the almost simple exceptional groups with $R(G)=6$, but it follows from [19, Corollary 3] that if $G_{0}=E_{6}(q)$ (respectively, $E_{7}(q)$), then the conjugate $5$-tuples containing copies of the $P_{1}$ or $P_{6}$ (respectively $P_{7}$) parabolic subgroup of $G$ are non-regular.
We now briefly discuss the methods we use to prove Theorem 2 for groups of Lie type. These are similar to those used in the proof of Cameron’s conjecture, but additional challenges arise due to the substantial weakening of our assumptions, which we will highlight below. Our main tool will be a generalised version of a powerful probabilistic method for studying base sizes, which was first introduced in [64] by Liebeck and Shalev.
Let $G$ be a finite group and let $\tau=(H_{1},\ldots,H_{k})$ be a tuple of core-free subgroups of $G$. We will write $X_{\tau}=G/H_{1}\times\cdots\times G/H_{k}$ and $\mathbb{P}(G,\tau)$ for the probability that a uniformly random point in $X_{\tau}$ is in a regular orbit of $G$. Then $\tau$ is regular if
@@
In general, it is very difficult to compute $Q(G,\tau)$ precisely. However, in most cases it is sufficient to work with the following upper bound (see [2]):
@@
where $x_{1},\ldots,x_{t}$ is a complete set of representatives of the prime order conjugacy classes in $G$ and
${\rm fpr}(x,G/H)=\frac{|C_{\Omega}(x)|}{|\Omega|}=\frac{|x^{G}\cap H|}{|x^{G}|}$
denotes the fixed point ratio of $x$ in the natural action of $G$ on the set $\Omega$ of cosets of $H$. That is, the proportion of fixed points of $x$ in this action. We refer the reader to Section 2.1 for more information.
Suppose that $G$ is an almost simple classical group with socle $G_{0}$ with natural module $V$ and let $\tau$ be a non-standard tuple of $G$. Here, our general approach when estimating $\widehat{Q}(G,\tau)$ is to use the following ‘zeta type function’
@@
where $t\in\mathbb{R}$ and $\mathscr{C}$ denotes the set ofr $G_{0}$-conjugacy classes of prime order elements of $G$. Burness obtains upper bounds on ${\rm fpr}(x,G/H)$ for a prime order element $x$ and a non-standard subgroup $H$ in terms of the size of the $G$-class of $x$ in [14, 15, 16, 17]. If ${\rm dim}V>10$, then using these allows us to bound $\widehat{Q}(G,\tau)$ by $\eta_{G}(t)$ for some specified $t\in(0,1)$. Our problem then reduces to showing that $\eta_{G}(t)<1$.
On the other hand, if ${\rm dim}V\leqslant 10$, then this general approach is no longer applicable. In some cases, we can work with a modified version of the general method, whereas a completely different approach is required in other cases. In particular, if ${\rm dim}V\leqslant 5$, using $\eta_{G}(t)$ will no longer be possible, so we take a more direct approach to estimate $\widehat{Q}(G,\tau)$. In these ‘low-dimensional’ cases, our proof heavily relies on a detailed analysis of fixed point ratios of classical groups in non-standard actions. More specifically, we make use of the description of the structure of the maximal subgroups of the low-dimensional classical groups given by Bray, Holt, and Roney-Dougal in [12], together with the detailed information on conjugacy classes of prime order elements presented in [22] by Burness and Giudici to improve the bounds given in [14, 15, 16, 17] sufficiently to force $\widehat{Q}(G,\tau)<1$. More information on this is given in Section 4. We believe that those results might be of independent interest, so we record them in Table 7.
Now suppose that $G$ is an exceptional group of Lie type over $\mathbb{F}_{q}$. Here, our main goal will be again to force $\widehat{Q}(G,\tau)<1$ for any $6$-tuple $\tau$ of core-free subgroups of $G$, but our approach when doing so is rather different to classical groups. Our proofs will again rely on fixed point ratio estimates, and our main reference for this will be [52]. In some cases, the uniform bounds given in [52, Theorems 1 and 2] will suffice for our purposes, but in most cases, a more detailed case-by-case analysis will be required. This can be achieved using the powerful machinery developed in [52] by Lawther, Liebeck and Seitz.
If $H$ is a non-parabolic subgroup of $G$, then upper bounds on ${\rm fpr}(x,G/H)$ for prime order elements $x\in G$ that are sufficient for our purposes can be extracted from [26]. Here, the authors use the techniques developed in [52] to derive those estimates. We believe that a record of such upper bounds could be of independent interest, so we record an upper bound on
$g(x,q)=\max\{{\rm fpr}(x,G/H)\,:\,H\leqslant G\text{ non-parabolic}\}$
for a prime order element $x\in G$ in Tables 13, 14, 15, 16, and 17 in Section 8 when $G_{0}\in\{F_{4}(q),E_{6}^{\epsilon}(q),E_{7}(q)\}$, where estimating $g(x,q)$ is a more laborious task.
The case where $H$ is a parabolic subgroup of $G$ requires a lot more work, particularly when $G_{0}\in\{F_{4}(q),E_{6}^{\epsilon}(q),E_{7}(q),E_{8}(q)\}$, in which case the bounds in [52, Theorems 1 and 2] are not good enough for our purposes. We note that if $H$ is a core-free maximal subgroup of $G$, then
${\rm fpr}(x,G/H)=\frac{\chi(x)}{\chi(1)},$
where $\chi(x)=1_{H}^{G}$ is the corresponding permutation character. Lawther, Liebeck and Seitz give an explicit formula for $\chi(x)$ in [52, 2.4] and [52, 3.2] for unipotent and semisimple elements respectively. Using these results together with information on conjugacy classes of unipotent and semisimple elements given in [60] and [3, 40, 41, 66] respectively, as well as data on the so-called Foulkes functions computed by Lübeck in [65], we are able to compute the precise value of ${\rm fpr}(x,G/H)$ for unipotent and semisimple elements.
For graph, field and graph-field automorphisms, we are again able to work as in [52] to obtain sufficient fixed point ratio estimates.
In the way we describe above, we are able to estimate
@@
for every prime order element $x\in G$, where $\mathcal{M}_{G}$ denotes the set of core-free maximal subgroups of $G$. This in turn allows us to obtain a uniform upper bound on $\widehat{Q}(G,\tau)$ that applies to any $6$-tuple $\tau$ of maximal subgroups of $G$.
The main challenge when the conjugacy assumption on our subgroup tuples is removed is that we often need to work with a very large number of possibilities of non-standard tuples, which forces us to adopt a uniform approach that allows us to handle all cases simultaneously. More specifically, our main strategy is to obtain a uniform upper bound on $\widehat{Q}(G,\tau)$ that applies to any choice of non-standard tuple $\tau$, which often means that for every choice of prime order element in $G$, we are bound to worst-case fixed point ratio estimates over all non-standard subgroups of our group $G$.
This means that we will often need to obtain stronger fixed point ratio estimates for our purposes. We state those in Theorem 4.1 for the classical groups. For exceptional groups, we record worst case fixed point ratio estimates over non-parabolic actions in Propositions 8.1, 8.2, 8.3, and 8.4. Our main source for obtaining those is inspection of [26] and [19], and in some instances we need to improve what currently is available. We give detailed information in Sections 2, 4 and 8.
@@
The probabilistic approach we adopt to prove Theorem 2 also allows us to establish some interesting asymptotic results. In particular, if $G$ is an almost simple group of Lie type and $\mathcal{P}$ denotes the set of non-standard subgroups of $G$, then we define
$\mathbb{P}(G,c)=\min\{\mathbb{P}(G,\tau)\,:\,\tau\in\mathcal{P}^{c}\}.$
We have the following theorem:
@@
We note that the conclusion to Theorem 4 also sheds light to the regularity of tuples of simple algebraic groups. We will explore this connection further in a follow-up paper.
@@
Organisation. This paper is organised as follows: In Section 2 we record some preliminary results on probabilistic and computational methods, and in Section 3, we present some preliminaries that only apply to classical groups. We discuss subgroup structure and conjugacy classes of classical groups, and we also give a precise definition of a subspace subgroup of a classical group (see Definition 3.4). The rest of the section is devoted to probabilistic methods, which is the key tool that we will be using.
@@
Section 4 is devoted to improving existing fixed point ratio estimates for low-dimensional classical groups. In Section 5 we prove some results relating to the ‘zeta-type’ function introduced in Definition 3.10, and Section 6 is devoted to the proof of Theorem 3.1, a refinement of Theorem 2 for classical groups.
@@
In Section 7 we record some preliminaries that only apply to exceptional groups. In particular, the section is divided into two subsections, where we briefly discuss subgroup structure and conjugacy classes respectively. In Section 8 we record fixed point ratio estimates for non-parabolic actions of exceptional groups, and then in Section 9, we proceed to discuss fixed point ratios for parabolic actions. Finally, Section 10 is devoted to the proof of Theorem 7.1, which allows us to conclude the proof of Theorem 2.
Notation. Our notation is fairly standard. If $n$ is a natural number, we will often write $[n]$ to denote the set $\{1,\ldots,n\}$. Moreover, we will also often write $\pi(n)$ to denote the set of prime divisors of a positive integer $n$ and $\pi_{a}(n)$ to denote the set of prime divisors $p$ of $n$ with $p\leqslant a$. In a similar fashion, we will denote the set of prime numbers at most $s$ by $\pi_{s}$ and the set of odd prime numbers at most $s$ by $\pi^{\prime}_{s}$.
@@
For a group $G$, we write $i_{s}(G)$ for the number of elements of order $s$ in $G$. For classical groups, we write ${\rm PSL}^{+}_{n}(q)={\rm PSL}_{n}(q)={\rm L}_{n}(q)$ and ${\rm PSL}^{-}_{n}(q)={\rm PSU}_{n}(q)={\rm U}_{n}(q)$. We use the same notation and terminology as [22] for field, graph, and graph-field automorphisms, and we follow the definition of [48] for the $\mathscr{C}_{i}$ and $\mathscr{S}$ collections of maximal subgroups of classical groups. We will also be using the standard notation $E_{6}^{\epsilon}(q)$ for the finite exceptional groups of type $E_{6}$, where $E_{6}^{+}(q)=E_{6}(q)$ and $E_{6}^{-}(q)={}^{2}E_{6}(q)$.
Finally, following the notation in [52], we will often write $x=s,u,\phi,\tau,\tau\phi$ to mean that $x$ is semisimple, unipotent, a field, graph, or graph-field automorphism of $G_{0}$ respectively.
We will introduce more technical terminology and notation, when and where appropriate in the paper.